5/8+7/8=12/8=1 1/2
Found 2 solutions by checkley79, carol1020:
Answer by checkley79(3341)

5/8+7/8=12/8=1.5 ANS.

Answer by carol1020(1)

5/8+7/8=12/8=1 1/2